A, B, C, D, and E are points on a circle. BFD and AFC are straight lines. DC = DF. Angle mass BAF = 24⁰. Angle mass AFE = x. EFD angle measure = 2x. Find the measure of the angle x.


1
Expert's answer
2021-05-28T10:00:19-0400



Now according to the diagram above , we can say that all three angles that is "( 24^0,x,2x)" lies on the same straight line .


According to the property of the straight line ,all angles lie on the same straight line at a point have the sum "= 180^0"


So accordingly:

"\Rightarrow" "x +2x+ 24 = 180 ^0"

"\Rightarrow 3x = 156^0"

"\Rightarrow x= 52^0"